OpenStudy (anonymous):

Anthony has a cup full of nickels and dimes that he would like to put in his savings account. When he takes the cup to the bank, the banker tells him that the 566 coins in the cup are worth $43.80. Which of the following systems of equations models the coins in Anthony’s cup? Let n represent nickels, and d represent dimes.

OpenStudy (anonymous):

The first thing I would do is change the $43 into cents so that equals..... if you don't know how to convert it let me know and I will help you figure out the first step

OpenStudy (texaschic101):

d = dimes n = nickels d + n = 566 --> d = 566 - n .10d + .05n = 43.80 Now just sub 566 - n in for d in the 2nd equation. Then solve for n. Once n is found, sub that back into the 1st equation and solve for d. If you want to check your answer, sub known values into the 2nd equation and if your answer comes out equal, then you have done it correctly....if not, then there is a mistake.